OpenStudy (anonymous):

the sum of the angles measures of a polygon with s sides is 2,520. Find s? a.)14 b.)16 c.)18 d.)15

OpenStudy (midhun.madhu1987):

Here it is "s" instead of "n"

OpenStudy (anonymous):

okay so you would....

OpenStudy (midhun.madhu1987):

Total angle of a polygon of n sides = (n-2) * 180 2520 = (n-2) * 180

OpenStudy (anonymous):

s=2520/180+2

OpenStudy (anonymous):

16 Thank you so much :)

OpenStudy (midhun.madhu1987):

yes